# Fano threefolds in positive characteristic III

### v3

@@ 7. Appendix: Computation of Chern classes
Let $Y$ be an $n$-dimensional smooth variety and let $\mathcal{E}$ be a locally free sheaf of rank $r$ on $Y$. Let $\pi\colon X=\mathbb{P}(\mathcal{E})\to Y$ be the projection induced by the $\mathbb{P}^{r-1}$-bundle structure. Then the following hold.
@@
    If $r=2$, then it holds that $\mathcal{O}_{X}(1)\cdot F=1$ for any fibre $F$ of $\pi:X\to Y$ over a closed point.
@@
where the second equality holds by $c_{2}(\mathcal{E})\cdot c_{1}(\mathcal{L})=0$ and the third one follows from $\pi^{*}Z\cdot\xi=\deg Z$ for any 00-cycle $Z\in\mathrm{A}^{2}(Y)$ (note that $\deg$ is dropped by abuse of notation). By $c_{1}(\mathcal{L})=c_{1}(\omega_{Y})+c_{1}(\det\mathcal{E})=K_{Y}+c_{1}(\mathcal{E})$, we obtain
@@
$\text{for}\qquad\mathcal{G}:=\omega_{Y}\otimes\det\mathcal{E}\otimes\mathcal{F}$
@@
where we used $c_{1}(\mathcal{E})\cdot c_{2}(\mathcal{E})=0$, which is guaranteed by $\dim Y=2$. By (7), (7.1.5) and (7.1.6),
@@
Since we have $\mathcal{O}_{X}(1)|_{F}=\mathcal{O}_{\mathbb{P}^{2}}(1)$ for every fibre $F\simeq\mathbb{P}^{2}$ of $\pi$, we have $\xi^{2}\cdot F=(\xi|_{F})^{2}=c_{1}(\mathcal{O}_{\mathbb{P}^{2}}(1))^{2}=1$, which implies $\pi^{*}Z\cdot\xi^{2}=\deg Z$ for any 00-cycle $Z\in\mathrm{A}^{2}(Y)$. It follows from (7) that
@@
The assertions (1) and (2) follow from [Har77, Ch. II, Exercise 8.5(b)] and and [Har77, II, Theorem 8.24(b),(c)], respectively. Let us show (3). By (2), we have $D\simeq\mathbb{P}(\mathcal{N}^{*}_{C/Y})$. We then get
@@
Let us show (4). Set $\pi:=f|_{D}\colon D=\mathbb{P}(\mathcal{N}^{*}_{C/Y})\to C$ and $\mathcal{L}:=(\omega_{C}\otimes\det\mathcal{N}^{*}_{C/Y})^{-1}$. Fix a divisor $\xi$ on $D$ such that $\mathcal{O}_{D}(\xi)=\mathcal{O}_{D}(1)$. Then
$(-K_{X})^{2}\cdot D=(\mathcal{O}_{X}(-K_{X})|_{D})^{2}.$
It holds that
@@ 8. Appendix: Description of split double covers
Assume $p=2$. Let $f:X\to Y$ be a double cover of smooth projective varieties. Set $\mathcal{L}:=(f_{*}\mathcal{O}_{X}/\mathcal{O}_{Y})^{-1}$, which is an invertible sheaf (Remark 2.2). Then $f$ is a spit double cover if and only if $f$ is a split $\alpha_{\mathcal{L},s}$-torsor associated to some $(\mathcal{L},s,t)$.
@@
Let $f:X\to Y$ be a split double cover of smooth projective varieties. Set $\mathcal{L}:=(f_{*}\mathcal{O}_{X}/\mathcal{O}_{Y})^{-1}$, which is an invertible sheaf (Remark 2.2). Let $\pi:Y\to Y^{\prime}$ be a morphism of smooth projective varieties such that $g_{*}\mathcal{O}_{Y}=\mathcal{O}_{Y^{\prime}}$ and $\mathcal{L}\simeq g^{*}\mathcal{L}^{\prime}$ for some invertible sheaf $\mathcal{L}^{\prime}$ on $Y^{\prime}$. Then there exist a split double cover $f^{\prime}:X^{\prime}\to Y^{\prime}$ and a morphism $\pi_{X}:X\to X^{\prime}$ such that the following diagram is cartesian:
@@
Assume $p=2$. By Lemma 8.3, there exist $s\in H^{0}(Y,\mathcal{L})$ and $t\in H^{0}(Y,\mathcal{L}^{\otimes 2})$ such that $f$ is the split $\alpha_{\mathcal{L},s}$-torsor associated to $(\mathcal{L},s,t)$. By $\pi_{*}\mathcal{O}_{Y}=\mathcal{O}_{Y^{\prime}}$ and $\mathcal{L}\simeq g^{*}\mathcal{L}^{\prime}$, we have the following corresponding elements:
@@ 9. Classification table of Fano threefolds with $\rho=2$
Let $X$ be a Fano threefold with $\rho(X)=2$. Then one of No. 2-1, …, No. 2-36 holds. In addition to (1)–(6) of Theorem 1.2, we use the follwoing notation and terminologies:
@@ 2-3 $8$ $D_{1}:(-K_{X})^{2}\cdot X_{t}=2$
$E_{1}:$ is a blowup of $V_{2}$ along an elliptic curve of degree $2$ which is a complete intersection of two members of $|-\frac{1}{2}K_{V_{2}}|$
@@ 2-14 $20$ $D_{1}:$ $(-K_{X})^{2}\cdot X_{t}=5$
2-15 $22$ $E_{1}:$ blowup of $\mathbb{P}^{3}$ along a curve of genus $3$ and degree $6$
@@ 2-21 $28$ $E_{1}:$ blowup of $Q$ along a rational curve of degree $4$
$28$ $E_{1}:$ blowup of $Q$ along a rational curve of degree $4$
@@ 2-28 $40$ $E_{1}:$ blowup of $\mathbb{P}^{3}$ along an elliptic curve of degree $3$
$E_{3}\,{\rm or}\,E_{4}$
@@ 2-36 $62$ $\mathbb{P}_{\mathbb{P}^{2}}(\mathcal{O}_{\mathbb{P}^{2}}\oplus\mathcal{O}_{\mathbb{P}^{2}}(2))$
(of Table LABEL:table-pic2) We use Notation 5.1. Almost all the parts follow from Section 5. It is enough to compute $\deg\Delta$ (type $C$) and $(-K_{X})^{2}\cdot X_{t}$ (type $D$).
If $f_{1}:X\to\mathbb{P}^{2}$ is of type $C$, then we have $H_{1}^{3}=0$ and $H_{1}^{2}\cdot H_{2}=(\frac{2}{\mu_{1}}\ell_{1})\cdot H_{2}=\frac{2}{\mu_{1}}$ (Lemma 5.3, Proposition 5.9), which imply the following:
@@
where (a) and (b) follow from Proposition 3.16 and Proposition 5.9, respectively. If $f_{1}:X\to\mathbb{P}^{1}$ is of type $D$, then we have $H_{1}^{3}=H_{1}^{2}\cdot H_{2}=0$, which imply the following:
$(-K_{X})^{2}\cdot X_{t}=(-K_{X})^{2}\cdot H_{1}=(\mu_{2}H_{1}+\mu_{1}H_{2})^{2}\cdot H_{1}=\mu_{1}^{2}H_{1}\cdot H_{2}^{2}.$
Therefore, it is enough to compute one of $(-K_{X})^{2}\cdot H_{1}$ and $H_{1}\cdot H_{2}^{2}$. If $R_{2}$ is type $D$, then we have $H_{2}^{2}\equiv 0$, and hence $H_{1}\cdot H_{2}^{2}=0$. If $R_{2}$ is of type $C$, then we get $H_{1}\cdot H_{2}^{2}=\frac{2}{\mu_{2}}$.
In what follows, we assume that $X$ is a Fano threefold of No. 2-8 in order to explain how to compute $\deg\Delta$ and $(-K_{X})^{2}\cdot H_{1}$ for the case when $R_{1}$ is of type $C$ and $R_{2}$ is of type $E$. It holds that $-K_{X}\sim H_{1}+H_{2}$ (Proposition 5.9). For $D:={\operatorname{Ex}}(f_{2})$, we have $K_{X}\sim f_{2}^{*}K_{Y}+D$ and $((-K_{X})^{2}\cdot D,(-K_{X})\cdot D^{2},D^{3})=(2,-2,2)$ [TanII, Proposition 3.22]. By Lemma 5.25(1), we have $D\cdot\ell_{1}=1$ (note that $\zeta\equiv 2\ell_{1}$ for a fibre $\zeta$ of $f_{1}:X\to\mathbb{P}^{2}$). Hence we obtain $-f_{2}^{*}K_{Y}\cdot\ell_{1}=-K_{X}\cdot\ell_{1}+D\cdot\ell_{1}=2$. By $H_{2}\cdot\ell_{1}=1$ (Proposition 5.9), we get $-f_{2}^{*}K_{Y}\equiv 2H_{2}$. Then
@@
We close this paper with the following proposition. Although this result is not used in this paper, we will need it in the classification of Fano threefolds with $\rho\geq 3$.
Proposition 9.1.
  1. (1)
  2. (2)
Proof.
We prove (1) and (2) simultaneously. If $X$ is 2-15 (resp. 2-23), then we set $Y:=\mathbb{P}^{3}$ (resp. $Y:=Q$) and $s:=4$ (resp. $s:=3$). Then we have a blowup $f:X\to Y$ and $Y$ is a Fano threefold of index $s$. We have the contraction $g:X\to Z$ of the other extremal ray, which is of type $E_{3}$ or $E_{4}$. In particular, $T:={\operatorname{Ex}}(g)$ is a (possibly singular) quadric surface. Set $S:=f(T)$. It is enough to show that $B\subset S$ and the induced composite morphism
[…]
is an isomorphism. Set $H_{f}$ (resp. $H_{g}$) to be the pullback on $X$ of the ample generator of $Y$ (resp. $Z$). It holds that $-K_{X}\sim H_{f}+H_{g}$ (Proposition 5.9). Let $r$ be the positive integer satisfying $-g^{*}K_{Y}\sim rH_{g}$.
Step 1.
$B\subset S$.
Proof of Step 1.
$00<S\cdot C=f^{*}S\cdot C_{X}=T\cdot C_{X}<$ 0 .
This completes the proof of Step 1. ∎
Step 2.
$T\cdot\zeta=r-1$ for a one-dimensional fibre $\zeta$ of $f:X\to Y$.
Proof of Step 2.
We have that
$-K_{X}\sim H_{f}+H_{g}\equiv\frac{-K_{X}+E}{s}+\frac{-K_{X}+T}{r},$
which implies
$-(sr-r-s)K_{X}\sim rE+sT.$
Step 3.
It holds that $r=2$ and $T\cdot\zeta=1$.
Proof of Step 3.
By [TanII, Proposition 3.22], we get
Step 4.
$f_{T}:T\to S$ is an isomorphism.
Proof of Step 4.
Let us show that $S$ is normal. Consider an exact sequence
$0\to\mathcal{O}_{X}(-T)\to\mathcal{O}_{X}\to\mathcal{O}_{T}\to 0$
By $(-T-K_{X})\cdot\zeta=-1+1=0$ (Step 3), we get $R^{1}f_{*}\mathcal{O}_{X}(-T)=0$ [Tan15, Theorem 0.5]. Then $\mathcal{O}_{S}=f_{*}\mathcal{O}_{T}$, i.e., $S$ is normal.
Note that any curve $\zeta$ on $T$ is not contracted by $f$, as otherwise we would get a contradiction: $\zeta\in R_{1}\cap R_{2}=\{0\}$. Hence $T\neq{\operatorname{Ex}}(f)$, which implies that $f:T\to S$ is birational. Since $T$ is a (possibly singular) quadric surface, $f|_{T}:T\to S$ is an isomorphism. This completes the proof of Step 4. ∎
Step 1 and Step 4 complete the proof of Proposition 9.1. ∎

### v4

@@ 7. Appendix: Computation of Chern classes
Let $Y$ be an $n$-dimensional smooth projective variety and let $\mathcal{E}$ be a locally free sheaf of rank $r$ on $Y$. Let $\pi\colon X=\mathbb{P}(\mathcal{E})\to Y$ be the projection induced by the $\mathbb{P}^{r-1}$-bundle structure. Then the following hold.
@@
    If $r=2$, then it holds that $\mathcal{O}_{X}(1)\cdot F=1$ for every fibre $F$ of $\pi:X\to Y$ over a closed point.
@@
where the second equality holds by $c_{2}(\mathcal{E})\cdot c_{1}(\mathcal{L})=0$ and the third one follows from $\pi^{*}Z\cdot\xi=\deg Z$ for every 00-cycle $Z\in\mathrm{A}^{2}(Y)$ (note that $\deg$ is dropped by abuse of notation). By $c_{1}(\mathcal{L})=c_{1}(\omega_{Y})+c_{1}(\det\mathcal{E})=K_{Y}+c_{1}(\mathcal{E})$, we obtain
@@
$\text{for}\qquad\mathcal{G}:=\omega_{Y}\otimes\det\mathcal{E}\otimes\mathcal{F}.$
@@
where we used $c_{1}(\mathcal{E})\cdot c_{2}(\mathcal{E})=0$, which is guaranteed by $\dim Y=2$. By (7), (7.1.5) and (7.1.6), we get
@@
Since we have $\mathcal{O}_{X}(1)|_{F}=\mathcal{O}_{\mathbb{P}^{2}}(1)$ for every fibre $F\simeq\mathbb{P}^{2}$ of $\pi$, it holds that $\xi^{2}\cdot F=(\xi|_{F})^{2}=c_{1}(\mathcal{O}_{\mathbb{P}^{2}}(1))^{2}=1$, which implies $\pi^{*}Z\cdot\xi^{2}=\deg Z$ for every 00-cycle $Z\in\mathrm{A}^{2}(Y)$. It follows from (7) that
@@
The assertions (1) and (2) follow from [Har77, Ch. II, Exercise 8.5(b)] and and [Har77, Ch. II, Theorem 8.24(b)(c)], respectively. Let us show (3). By (2), we have $D\simeq\mathbb{P}(\mathcal{N}^{*}_{C/Y})$. We then get
@@
Let us show (4). Set $\pi:=f|_{D}\colon D=\mathbb{P}(\mathcal{N}^{*}_{C/Y})\to C$ and $\mathcal{L}:=(\omega_{C}\otimes\det\mathcal{N}^{*}_{C/Y})^{-1}$. Fix a divisor $\xi$ on $D$ such that $\mathcal{O}_{D}(\xi)=\mathcal{O}_{D}(1)$. It holds that
@@ 8. Appendix: Description of split double covers
Assume $p=2$. Let $f:X\to Y$ be a double cover of smooth projective varieties. Set $\mathcal{L}:=(f_{*}\mathcal{O}_{X}/\mathcal{O}_{Y})^{-1}$, which is an invertible sheaf (Remark 2.2). Then $f$ is a split double cover if and only if $f$ is a split $\alpha_{\mathcal{L},s}$-torsor associated to some $(\mathcal{L},s,t)$.
@@
Let $f:X\to Y$ be a split double cover of smooth projective varieties. Set $\mathcal{L}:=(f_{*}\mathcal{O}_{X}/\mathcal{O}_{Y})^{-1}$, which is an invertible sheaf (Remark 2.2). Let $\pi:Y\to Y^{\prime}$ be a morphism of smooth projective varieties such that ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi}_{*}\mathcal{O}_{Y}=\mathcal{O}_{Y^{\prime}}$ and $\mathcal{L}\simeq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi}^{*}\mathcal{L}^{\prime}$ for some invertible sheaf $\mathcal{L}^{\prime}$ on $Y^{\prime}$. Then there exist a split double cover $f^{\prime}:X^{\prime}\to Y^{\prime}$ and a morphism $\pi_{X}:X\to X^{\prime}$ such that the following diagram is cartesian:
@@
Assume $p=2$. By Lemma 8.3, there exist $s\in H^{0}(Y,\mathcal{L})$ and $t\in H^{0}(Y,\mathcal{L}^{\otimes 2})$ such that $f$ is the split $\alpha_{\mathcal{L},s}$-torsor associated to $(\mathcal{L},s,t)$. By $\pi_{*}\mathcal{O}_{Y}=\mathcal{O}_{Y^{\prime}}$ and $\mathcal{L}\simeq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi}^{*}\mathcal{L}^{\prime}$, we have the following corresponding elements:
@@ 9. Classification table of Fano threefolds with $\rho=2$
Definition 9.1.
We say that $X$ is a Fano threefold of No.  2-xx if $X$ is a Fano threefold with $\rho(X)=2$ such that $(-K_{X})^{3}$ and the types of the extremal rays are as in Table LABEL:table-pic2 below. For example, No. 2-1 and No. 2-15 are defined as follows.
  • $X$ is a Fano threefold of No. 2-1 if $X$ a Fano threefold such that $\rho(X)=2$, $(-K_{X})^{3}=4$, one of the extremal rays is of type $D_{1}$, and the other is of type $E_{1}$.
  • $X$ is a Fano threefold of No. 2-15 if $X$ a Fano threefold such that $\rho(X)=2$, $(-K_{X})^{3}=22$, one of the extremal rays is of type $E_{1}$, and the other is of type $E_{3}$ or $E_{4}$.
Let $X$ be a Fano threefold with $\rho(X)=2$. Then one of No. 2-1, …, No. 2-36 holds. In Table LABEL:table-pic2, we use the following notation and terminologies.
  1. (1)
    We say that $f:X\to Y$ is a split double cover if $f$ is a finite surjective morphism of projective normal varieites such that $\mathcal{O}_{Y}\to f_{*}\mathcal{O}_{X}$ splits as an $\mathcal{O}_{Y}$-module homomorphism and the induced field extension $K(X)\supset K(Y)$ is of degree two. For a split double cover $f:X\to Y$, we set $\mathcal{L}:=(f_{*}\mathcal{O}_{X}/\mathcal{O}_{Y})^{-1}$, which is an invertible sheaf on $Y$ (Remark 2.2).
  2. (2)
    $Q$ is the smooth quadric hypersurace in $\mathbb{P}^{4}$.
  3. (3)
    $W$ is a smooth prime divisor on $\mathbb{P}^{2}\times\mathbb{P}^{2}$ of bidegree $(1,1)$. Note that such a threefold is unique up to isomorphisms (Lemma 5.16).
  4. (4)
  5. (5)
@@ 2-3 $8$ $D_{1}:(-K_{X})^{2}\cdot X_{t}=2$
$E_{1}:$ blowup of $V_{2}$ along an elliptic curve of degree $2$ which is a complete intersection of two members of $|-\frac{1}{2}K_{V_{2}}|$
@@ 2-14 $20$ $D_{1}:$ $(-K_{X})^{2}\cdot X_{t}=5$
[…]
@@ 2-21 $28$ $E_{1}:$ blowup of $Q$ along a rational curve of degree $4$
$E_{1}:$ blowup of $Q$ along a rational curve of degree $4$
@@ 2-28 $40$ $E_{1}:$ blowup of $\mathbb{P}^{3}$ along an elliptic curve of degree $3$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}E_{5}}$
@@ 2-36 $62$ $\mathbb{P}_{\mathbb{P}^{2}}(\mathcal{O}_{\mathbb{P}^{2}}\oplus\mathcal{O}_{\mathbb{P}^{2}}(2))$
(of Table LABEL:table-pic2) We use Notation 5.1. Almost all the parts follow from Section 5. It is enough to compute
  1. (D)
    $(-K_{X})^{2}\cdot X_{t}$ (type $D$), and
  2. (C)
    $\deg\Delta$ (type $C$).
(D) Assume that $f_{1}:X\to\mathbb{P}^{1}$ is of type $D$ and let us see how to compute $(-K_{X})^{2}\cdot X_{t}$. Recall that $f_{2}$ is of type $E_{1}$ or $C$ (Theorem 4.17(2)). If $f_{2}$ is of type $E_{1}$, then $(-K_{X})^{2}\cdot X_{t}$ has been already comuted in Proposition 5.41. Assume that $f_{2}$ is of type $C$. In this case, we have $H_{1}^{3}=H_{1}^{2}\cdot H_{2}=0$ and $H_{2}^{2}\equiv(2/\mu_{2})\ell_{2}$ (Corollary 4.5), which imply the following:
$(-K_{X})^{2}\cdot X_{t}=(\mu_{2}H_{1}+\mu_{1}H_{2})^{2}\cdot H_{1}=\mu_{1}^{2}H_{1}\cdot H_{2}^{2}=\frac{2\mu_{1}^{2}}{\mu_{2}}H_{1}\cdot\ell_{2}=\frac{2\mu_{1}^{2}}{\mu_{2}},$
where the last equality holds by Proposition 5.9(2).
(C) Assume that $f_{1}:X\to\mathbb{P}^{2}$ is of type $C$ and let us see how to compute $\deg\Delta$ for $\Delta:=\Delta_{f_{1}}$. We have $H_{1}^{3}=0$ and $H_{1}^{2}\cdot H_{2}=(\frac{2}{\mu_{1}}\ell_{1})\cdot H_{2}=\frac{2}{\mu_{1}}$ (Lemma 5.3, Proposition 5.9), which imply the following:
@@
where (a) and (b) follow from Proposition 3.16 and Proposition 5.9, respectively. If $f_{2}$ is of type $C$ (resp. $D$), then we get $H_{1}\cdot H_{2}^{2}=H_{1}\cdot(\frac{2}{\mu_{2}}\ell_{2})=\frac{2}{\mu_{2}}$ (resp. $H_{1}\cdot H_{2}^{2}=0$).
What is remaining is the case when $f_{2}$ is of type $E$. If $f_{2}$ is of type $E_{1}$, then $\deg\Delta$ has already been computed in Proposition 5.36. Hence we may assume that $f_{2}$ is of type $E_{2},E_{3},E_{4}$, or $E_{5}$. By Theorem 5.33, the problem is reduced to the case when $X$ is of No. 2-8. It holds that $-K_{X}\sim H_{1}+H_{2}$ (Proposition 5.9). For $D:={\operatorname{Ex}}(f_{2})$, we have $K_{X}\sim f_{2}^{*}K_{Y}+D$ and $((-K_{X})^{2}\cdot D,(-K_{X})\cdot D^{2},D^{3})=(2,-2,2)$ [TanII, Proposition 3.22]. By Lemma 5.24(1), we have $D\cdot\ell_{1}=1$ (note that $\zeta\equiv 2\ell_{1}$ for a fibre $\zeta$ of $f_{1}:X\to\mathbb{P}^{2}$). Hence we obtain $-f_{2}^{*}K_{Y}\cdot\ell_{1}=-K_{X}\cdot\ell_{1}+D\cdot\ell_{1}=2$. By $H_{2}\cdot\ell_{1}=1$ (Proposition 5.9), we get $-f_{2}^{*}K_{Y}\equiv 2H_{2}$. Then